OpenStudy (anonymous):

The length of a rectangle is 48 in and is dilated to 12 in. What is the scale factor? A. one-fourth B. -4 C. 4 D. negative one-fourth Is it four?

OpenStudy (dan815):

12*4=48 12=48/4 12=1/4 * 48